OpenStudy (anonymous):

i forgot how to solve these, can someone teach me again? 2 2/5x 3 3/8

OpenStudy (mertsj):

Change both mixed numbers to improper fractions first.

OpenStudy (mertsj):

What would that be?

OpenStudy (anonymous):

12/5 and 27/8

OpenStudy (mertsj):

Very good.

OpenStudy (anonymous):

yay!

OpenStudy (mertsj):

\[\frac{12}{5}\times\frac{27}{8}\]

OpenStudy (mertsj):

Now cancel a 4 from the 12 ands 8

OpenStudy (mertsj):

\[\frac{3}{5}\times\frac{27}{2}\]

OpenStudy (mertsj):

Now multiply numerators and denominators and change back to a mixed number.

OpenStudy (anonymous):

wait im confused on the canceling part

OpenStudy (mertsj):

\[\frac{3}{5}\times\frac{27}{2}=\frac{81}{10}=8 \frac{1}{10}\]

OpenStudy (mertsj):